• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Sulkosky

    Log entry time 16:05:21 on June21,2001

    Entry number 64035

    keyword=Shift Summary

    8:00 Shift began waiting for beam, no estimate on how long. Present configuration is Holding Field at 0 degrees. Both half-wave plates OUT.
    Pol = 39%

    8:30 MCC called. They are giving us pulsed beam. Move target to BeO to check spot and raster.

    8:55 MCC called. They said they are still having some problems with the helium cryo and RF. No estimated time was available.

    9:30 EPR polarization: 38.1 %

    10:40 MCC called. They are giving us CW beam to do some tests.

    10:50 MCC done with test. 3 muA of CW on BeO target to check spot and raster.
    Position looks good. Raster looks good.

    10:55 Moved target to Pol 3He
    11:00 Taking data on Pol 3He at 0 degrees both half-waveplates OUT. grad coils at (L-3.00,T6.00)

    11:35 MCC called. They said they checked the problem out and are ready to give us CW beam again.

    14:55 Beam to hall A stopped for magnet work. Estimate about 45 minutes.

    15:00 EPR Measurement. Shape didn't look good. Called Xiaochao to come.

    15:45 Xiaochao began an EPR measurement.

    Run summary:

    Runs 1692 and 1693 cosmic runs.

    1695/20695 11:01-11:24 Production. Stopped Box supply went off. 15 minutes to investigate.

    1696/20696 11:40-12:34 Production. Stopped Box tripped off again. 10 minutes to fix.

    1697/20697 12:43-13:45 Production. some beam trips.
    1698/20698 13:48-14:53 Production. Charge asymmetry was high about 450 ppm. I tried to time stamp a value but it went higher. After talking to Bob Michaels and MCC, the slope of the rotatable waveplate was wrong. MCC was to correct the problem.